OpenStudy (anonymous):

add: (-4 + n - 2n2) + 2 - 5n - 6n2

OpenStudy (anonymous):

((((−4+n)−2n(2))+2)−5n)−6n(2) =(−4)+n+(−4n)+2+(−5n)+(−12n) =(−20n)+(−2) =−20n−2

OpenStudy (anonymous):

when adding variables like this, treat them like normal numbers but just deal with the coefficient differently.